UT Mar 29 2021 - Full D/E Night



Notes

  • Roll over science and engineering tasks from UT Mar 14.
  • 26 deg elevation limit in place for first few hours

Night Plan

  • Twilight to 21:30hr
    • DONE LBTI DDT Sirius Observations - hits 26 deg limit at 21:28 hr
      • Conditions: Any
      • Time Needed: twilight until 21hr
    • MODS Followup of Myster Object 21hr-22hr - requires 26 deg elevation change - hits 26 deg at 21:12hr
      • Spec Std after 21:12 hr
      • if conditions too poor for LBTI or other technical issues- run this program - get whatever photons can be obtained
      • If LBTI finishes early, and there's time to get some data before target sets - switch to MODS.
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/ICRF3/MODS_Observations
      • OT File: ICRF3_LUCI_V15_LBTO_AZ.xml
      • LBTO: LD-2019B-006
      • 1" slit Dual Grating
  • Reconfigure 21:30hr to LBCs
  • DONE LBC TMS Testing - 21:50 hr - 00:50 hr
    • Any conditions acceptable
  • Reconfigure for LUCI-AO/ESM 00:50hr
  • DONE Activate AO Software
  • DONE OSCO LUCI-1 01:20-01:50hr
  • DONE IMBH Candidate - ESM Ks imaging - 01:50hr - 03:50 hr (if photometric - standards add time)
    • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/IMBH/LUCI-ESM
      • OT file: J1601_Optical_Followup_V3.xml
    • Take mixed mode - SL data used for characterization of the ESM performance
    • Conditions: Seeing < 1.2", very thin clouds, photometric preferred
    • If Photometric - observe photometric standard stars before and after - 1hr - Grand Total - 2hrs
      • GSPC P330-E - zJHKKs - observe all filters to add more data for ZPs
      • adds about 1 hr of overheads (30 min before target, 30 min after target)
  • LUCI-1 AO check for pupil "shadows" - Bin 1 bright target (something like Titan would suffice)
    • Executed after IMBH target is complete
  • ESM Characterization 03:50hr-twilight
    • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • Hoag's Object - highest priority - on-axis AO Ref star
    • M92 - if time
  • Poor weather/conditions for 01:50 hr onwards
    • ESM observations
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
      • M51 - Ks imaging - 1hr
      • M92 - all filters - 1 hr
    • Photometric Standards - only if clear and seeing is poor
    • CO Mergers LD-2019B-005
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
      • UGC 9829 - if time after ESM M92 (or if can't close loop with ESM)
      • NGC 6598 - if time after ESM M92 (or if can't close the loop)
  • If M3 SX fails to move to LUCI-1
    • Switch to MODS
    • MODS Field Correction check for IT 8315 - Engineering Task run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
      • Seeing < 1.4" - thin clouds ok
      • do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
    • Spectrophotometric Standards - any seeing must be clear
    • Photometric Standards - any seeing must be clear
    • Local Mergers MODS Observations LD-2019B-005
      • any conditions
      • NGC 6598
    • RLQs/RQQs LD-2019B-004
      • any coditions
      • PG 1307+085
      • PG 1416-129
      • HB89 1635+119

Log available here:

https://docs.google.com/document/d/1dbLtKv-6Ta0DvbHUzd4AhYWWYKDTu2vYSaTTPXS86hw/edit?usp=sharing

-- BarryRothberg - 23 Mar 2021
Topic revision: r10 - 07 Apr 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback